OpenStudy (anonymous):

Choose the correct simplification of x to the 2nd power over x to the 5th power. 1 over x to the 3rd power x3 x7 1 over the x to the 7th power

OpenStudy (anonymous):

so its B?

OpenStudy (anonymous):

\[\frac{ x^2}{x^5} \frac{ 1}{ x^2-^5} = x^3\]

OpenStudy (anonymous):

5-2 = 3 so x^3

OpenStudy (anonymous):

B, x^3

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

m=2 n=5

OpenStudy (anonymous):

x^5-^2= x^3